OpenStudy (anonymous):

a state's high school athletic association separates its schools into two divisions. together div I and II have a total of 300 schools. there are about 6 times as a many schools in div I than div II. how many schools are in eatch div? just devided by 300/6 for div II. and 300-div II for div I?? or is there any formula???

OpenStudy (compassionate):

Hint: Use a system of equations.

OpenStudy (compassionate):

Do you know what that is?

OpenStudy (anonymous):

a = div I b= div II 6a+b=300????

OpenStudy (compassionate):

No, you're method is good, but let Div: 1 = x, and Div: 2 = y That is, x + y = 300

OpenStudy (anonymous):

6 times as a many schools in div I than div II tho..???

OpenStudy (compassionate):

...you need to define your variables ! let x = # of schools in division I let y = # of schools in division II then, x + y = 300 x = 6y

OpenStudy (compassionate):

The first part says Div: 1, and Div: 2 = 300 that is, x + y = 300 The second part says, "there are about 6 times as a many schools in div I than div II." Thus, x = 6y

OpenStudy (compassionate):

However, it says, "About 6 times." x + y = 300 x ≈ 6y

OpenStudy (compassionate):

^ Can you take it from here?

OpenStudy (anonymous):

thanks ;]

OpenStudy (compassionate):

No problem.